On a trip Joes drove 25 miles more than half the number of miles Rita drove. Together they drove 340 miles. How many miles did Joes drive?
Answer by mananth(16949)

let Rita drive x miles
1/2 x = x/2
joes drove x/2 + 25
x + x/2 +25 = 340
1.5x +25 = 340
1.5x = 340-25
1.5x=315
x=315/1.5
x=210 miles
Rita drove 210 miles
Joes drove 130 miles

Ananth